Circular No. 6716

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997dc IN NGC 7678 W.-d. Li, Beijing Astronomical Observatory, reports that a low-resolu-

tion spectrogram (20 nm/mm, range 380–830 nm) of SN 1997dc was obtained with the 2.16-m telescope at Xinglong station by Li, W. Liu, Q.-y. Qiao, and A. Esamdin. The spectrum shows broad emission and absorption lines typical of a type-Ic supernova at maximum light. The whole spectrum is very similar to that of SN 1994I at maximum light (Filippenko et al. 1995, Ap.J. 450, L11). Strong spectral lines include the Ca II infrared triplet and H and K lines; O I at 774 nm; Na I D/He I at 580 nm; Si II at 630 nm; and Fe II lines at 445, 490, and 508 nm. There is no distinct evidence of H Balmer lines.

V1333 AQUILAE C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence,

write: “CCD monitoring of the optical counterpart of the recurrent lowmass x-ray transient Aql X-1 with the OHP 1.2-m telescope on Aug. 6.9, 7.9, 8.9, 9.9, and 10.9 UT shows that the object is hovering around V = 17.5, indicating that the current outburst has probably reached a maximum level typical of previous such mild events (see IAUC 6638). Recent publicly available XTE data (up to Aug. 11.29) show that the x-ray flux reported on IAUC 6714 indeed appears to have leveled off since Aug. 8.5.”

SUPERNOVAE 1997bp AND 1997cx S. Moretti and S. Tomaselli, Forlí, Italy, report their unfiltered CCD

magnitudes of recent supernovae (GSC comparison stars): SN 1997bp in NGC 4680, May 27.92 UT, 16.1; SN 1997cx in NGC 3057, July 28.96, 15.3.

COMET 78P/GEHRELS 2 Total CCD magnitudes: Aug. 9.07 UT, 12.6 (A. Galád and A. Pravda,

Modra, Slovakia, 0.6-m f/5.5 reflector); 12.08, 13.1 (H. Mikuž, Crni Vrh Observatory, Slovenia, 0.36-m f/6.7 reflector + V filter).

CORRIGENDUM On IAUC 6715, the last two magnitude estimates attributed to comet

C/1997 J1 (Mueller) belong to comet C/1996 J1 (Evans-Drinkwater).

1997 August 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green